Lacrosse. The perimeter of an NCAA men’s lacrosse field is 340 yd. The length is 50 yd longer than the width. Find the dimensions.
Answer by ikleyn(53751)

.
L + W + L + W = 340   yards   (the perimeter)

2L + 2W = 340


L + W = 340/2 = 170.


Substitute L = W + 50.


(W + 50) + W = 170


2W = 170 - 50 = 120


W = 120/2 = 60.


Answer.  The dimensions are  60 yd (width)  and  60 + 50 = 110 yd (length).

Solved.